OpenStudy (anonymous):

use the substitution method to solve the given system of equations.write your answer as an ordered pair (x,y). 3x-4y=13 and x-2y=9

OpenStudy (smarty):

x=2y+9 Substitute that equation into the first one so: 3(2y+9)-4y=13 6y+27-4y=13 Combine like terms 2y=-14 y=-7 Then plug -7 for y in any of the equations x-2(-7)=9 x+14=9 x=-5 (-5,-7) Plug those numbers to check if it is done correctly.